AM4 -Esercitazione 6 A.A.2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 30 ottobre 2003

Esercizio 1

Sia Γ il triangolo di vertici (1, 0, 0), (0, 1, 0), (0, 0, 1). Calcolare ∫_{Γ} yz\sqrt{1 − x} dσ.

Esercizio 2

Calcolare

∫_{γ} x^{2} ln z ds

dove γ := (x = e^{t} cos t, y = e^{t} sin t, z = e^{t}) con 0 ≤ t ≤ T.

Esercizio 3

Calcolare i seguenti integrali indefiniti:

a) ∫ e^{t} ln \frac{1}{\sqrt{1 − e^{−t}}} dt;

b) ∫ \frac{ln^{2} x}{x^{2}} dx